ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: Bamba, Kanté Siaka, Universite de Cocody, UFR Mathematiques 22 B. P. 1079 Abidjan 22 On quadratic modules I build quad a generalized category of quadratic modules with direct (lim) and inverse (lim) limits. We show that the functor lim in the category of direct systems, preserves short exact sequences. In particular lim is exact. I thank the Department of Mathematics of the University of Illinois at Urbana-Champaign for its hospitality.